# from early-time inflation to late-time acceleration

### v1

@@ I Introduction
In this paper, we show that the symmetric teleparallel theory does not exclude the Riemann curvature constructed from the standard Levi-Civita connection. This idea is not new, of course. In Gadbail:2023mvu , a model called $f(Q,C)$ gravity is proposed, where $C$ is the difference between $Q$ and the scalar curvature $\tilde{R}$ in Einstein’s gravity, $C=\tilde{R}-Q$. We should note that $C$ is a total derivative. The action of $f(Q,C)$ gravity is given by a function $f(Q,C)$ of $Q$ and $C$. Even if we consider the $f(Q,\tilde{R})$ gravity, whose action is given by a function $f(Q,\tilde{R})$, because $f(Q,\tilde{R})=f(Q,Q+C)$, the model can be regarded as the $f(Q,C)$ gravity. One reason why the symmetric teleparallel theory can include the curvature in Einstein’s gravity is that the conservation law is usually given in terms of the covariant derivative defined by the Levi-Civita connection. The commutator of two covariant derivatives inevitably induces the curvature in Einstein’s gravity. In this paper, we propose to consider $f\left(Q,\mathcal{G}\right)$ gravity, where $\mathcal{G}$ is the Gauss-Bonnet invariant. We apply this model to the cosmology and give an explicit formulation for the cosmological reconstruction. It is proposed a systematic way to construct such a theory which realises any given geometry.
@@ III $Q\mathcal{G}$ gravity
Therefore the covariant derivative given by only the Levi-Civita connection ${\tilde{\Gamma}^{\sigma}}_{\mu\nu}$ appears even in the $f(Q)$ gravity. The commutator of the Levi-Civita covariant derivatives gives the Riemann curvature in the Einstein gravity. The curvatures do not, of course, conflict with the general covariance. Therefore one may consider the model of $f(Q)$ gravity including the curvatures in the Einstein gravity, where the Lagrangian density is given by the function of $Q$ and the scalar curvature in the Einstein gravity $\tilde{R}$, that is, $f(Q,\tilde{R})$, or even the Gauss-Bonnet invariant $\mathcal{G}$, $f(Q,\tilde{R},\mathcal{G})$, where
@@
In the following, we consider $Q\mathcal{G}$ gravity, where the Lagrangian density is a function of $Q$ and $\mathcal{G}$, $f(Q,\mathcal{G})$:
$S_{Q\mathcal{G}}=\int d^{4}x\sqrt{-g}f\,.$ (20)
@@ IV Ghost problem
Let us remark that the Einstein-$f\left(\mathcal{G}\right)$ gravity includes ghosts in general. Nevertheless, some ghost-free models were proposed in Nojiri:2018ouv . In $f(Q,\tilde{R})$ or $\mathcal{G}$, $f(Q,\tilde{R},\mathcal{G})$, there could appear the ghost. First, we briefly review the ghost problem in the Einstein-$f\left(\mathcal{G}\right)$ gravity.
The action of the Einstein-$f\left(\mathcal{G}\right)$ gravity Nojiri:2005jg ; Cognola:2006eg ; Leith:2007bu ; Li:2007jm ; Kofinas:2014owa ; Zhou:2009cy is given by
$S=\int d^{4}x\sqrt{-g}\left(\frac{1}{2\kappa^{2}}R+f(\mathcal{G})+\mathcal{L}_{\mathrm{matter}}\right)\,.$ (24)
@@
$S=\int d^{4}x\sqrt{-g}\left(\frac{1}{2\kappa^{2}}R-\frac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi+h\left(\chi\right)\mathcal{G}-V\left(\chi\right)+\mathcal{L}_{\mathrm{matter}}\right)\,.$ (25)
@@
$S=\int d^{4}x\sqrt{-g}\left(\frac{1}{2\kappa^{2}}R+h\left(\chi\right)\mathcal{G}-V\left(\chi\right)+\mathcal{L}_{\mathrm{matter}}\right)\,.$ (26)
@@
$S=\int d^{4}x\sqrt{-g}\left\{\frac{1}{2\kappa^{2}}R+\lambda\left(\frac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi+\frac{\mu^{4}}{2}\right)-\frac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi+h\left(\chi\right)\mathcal{G}-V\left(\chi\right)+\mathcal{L}_{\mathrm{matter}}\right\}\,.$ (29)
@@
V FLRW spacetime in $f$ gravity and cosmological reconstruction
@@ VI.4 Autonomous dynamical system
Here we have used the relation $\frac{d}{dt}=H\frac{d}{dN}$ between the derivatives with respect to the cosmological time $t$ and $e$-folding number $N$. For the autonomous dynamical system, it is often used dimensionless quantities but because we have not specified $f\left(Q,\mathcal{G}\right)$, we use the dimensional quantities like $Q$ and $\mathcal{G}$. The fixed point defined by $\frac{d\mathcal{G}}{dN}=\frac{dQ}{dN}=0$ corresponds to de Sitter spacetime, where the Hubble rate $H$ is a constant $H=H_{0}$. Then the conditions for the fixed point are,
@@ VI.5 Gravitational wave
Note that $Q$ and $\mathcal{G}$ are invariant under the variation $g_{\mu\nu}\to g_{\mu\nu}+h_{\mu\nu}$ with (98). We are also now interested in the propagation speed of the gravitational wave and therefore we only investigate the terms including the second order derivative of $h_{\mu\nu}$. Then Eq. (III) tells
@@
In the FLRW spacetime, where $Q$ and $\mathcal{G}$ only depends on the cosmological time $t$, the condition (100) can be satisfied is $f_{\mathcal{G}}$ is a linear function of $t$, $f_{\mathcal{G}}=f_{0}+f_{1}t$. Here $f_{0}$ and $f_{1}$ are constants. The condition (100) cannot be, however, satisfied for the static and spherically symmetric spacetime as in the case of the Einstein-Gauss-Bonnet theories Nojiri:2023jtf .
@@ VII $QR\mathcal{G}$ gravity
As we mentioned, $f(Q,\tilde{R})$ gravity is identical to $f(Q,C)$ gravity proposed in Gadbail:2023mvu . So far, we have discussed $f$ gravity in detail. In this section, we briefly discuss a generalized example, that is, $f(Q,\tilde{R},\mathcal{G})$ gravity for completeness,
$S_{\mathrm{QR}\mathcal{G}}=\int d^{4}x\sqrt{-g}f\left(Q,R,\mathcal{G}\right)\,.$ (101)
@@
$\,-\tilde{R}_{\mu\nu}f_{R}+\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}f_{R}-g_{\mu\nu}\tilde{\nabla}^{2}f_{R}-2f_{\mathcal{G}}{\tilde{R}}{\tilde{R}}_{\mu\nu}+4f_{\mathcal{G}}{\tilde{R}}_{\mu\rho}{\tilde{R}}_{\nu}^{\ \rho}-2f_{\mathcal{G}}{\tilde{R}}_{\mu}^{\ \rho\sigma\tau}{\tilde{R}}_{\nu\rho\sigma\tau}$
@@
$S_{\mathrm{general}}=\int d^{4}x\sqrt{-g}f_{\mathrm{general}}\left(Q,g_{\mu\nu},R_{\mu\nu\rho\sigma}\right)\,.$ (103)
Here $g_{\mu\nu}$ expresses the metric which is not included in $Q$ and $R_{\mu\nu\rho\sigma}$ in the arguments. The equation corresponding to (VII) can be easily found by using the following formula of the variation with respect to the metric,
@@
Of course, $f\left(Q,\mathcal{G}\right)$ gravity and $f\left(Q,R,\mathcal{G}\right)$ gravity are special cases of the general model (103). For example, $f\left(Q,R,\mathcal{G}\right)$ corresponds to the following $f_{\mathrm{general}}\left(Q,g_{\mu\nu},R_{\mu\nu\rho\sigma}\right)$,
$f_{\mathrm{general}}$ $\,\left(Q,g_{\mu\nu},R_{\mu\nu\rho\sigma}\right)$
$=$ $\,f\left(Q,R=g^{\mu\nu}g^{\rho\sigma}R_{\mu\rho\nu\sigma},\right.$
@@
×$\qquad\qquad\left.\times R_{\alpha\beta\gamma\delta}R_{\mu\nu\rho\sigma}\right)\,.$ (105)

### v2

@@ I Introduction
In this paper, we show that the symmetric teleparallel theory does not exclude the Riemann curvature constructed from the standard Levi-Civita connection. This idea is not new, of course. A model called $f(Q,B)$ gravity in Capozziello:2023vne or $f(Q,C)$ gravity in Gadbail:2023mvu is proposed, where $B$ or $C$ is the difference between $Q$ and the scalar curvature $\tilde{R}$ in Einstein’s gravity, $B=Q-\tilde{R}$ or $C=\tilde{R}-Q$. For the application of $f(Q,B)$ gravity to cosmology, see Paliathanasis:2023kqs ; Paliathanasis:2024yea . We should note that $B$ or $C$ is a total derivative. The action of $f(Q,B)$ or $f(Q,C)$ gravity is given by a function $f(Q,B)$ or $f(Q,C)$ of $Q$ and $B$ or $C$. Even if we consider the $f(Q,\tilde{R})$ gravity, whose action is given by a function $f(Q,\tilde{R})$, because $f(Q,\tilde{R})=f(Q,Q-B)=f(Q,Q+C)$, the model can be regarded as the $f(Q,B)$ or $f(Q,C)$ gravity. One reason why the symmetric teleparallel theory can include the curvature in Einstein’s gravity is that the conservation law is usually given in terms of the covariant derivative defined by the Levi-Civita connection. The commutator of two covariant derivatives inevitably induces the curvature in Einstein’s gravity. In this paper, we propose to consider $f\left(Q,\mathcal{G}\right)$ gravity, where $\mathcal{G}$ is the Gauss-Bonnet invariant. We apply this model to the cosmology and give an explicit formulation for the cosmological reconstruction. It is proposed a systematic way to construct such a theory which realises any given geometry.
@@ III $Q\mathcal{G}$ gravity
Therefore the covariant derivative given by only the Levi-Civita connection ${\tilde{\Gamma}^{\sigma}}_{\mu\nu}$ appears even in the $f(Q)$ gravity. The commutator of the Levi-Civita covariant derivatives gives the Riemann curvature in the Einstein gravity. The curvatures do not, of course, conflict with the general covariance. Therefore one may consider the model of $f(Q)$ gravity including the curvatures in the Einstein gravity, where the Lagrangian density is given by the function of $Q$ and the scalar curvature in the Einstein gravity $\tilde{R}$, that is, $f(Q,\tilde{R})$, or even the Gauss-Bonnet invariant $\mathcal{G}$, $f(Q,\tilde{R},\mathcal{G})$, whose action is given by
$S_{\mathrm{QR}\mathcal{G}}=\int d^{4}x\sqrt{-g}f\left(Q,\tilde{R},\mathcal{G}\right)\,.$ (19)
where the Gauss-Bonnet invariant $\mathcal{G}$ is defined by
@@
Because $f(R)$ gravity includes the scalar mode, $f(Q,\tilde{R},\mathcal{G})$ could also include the scalar mode. For this reason, in the following, we consider a simpler model, that is, $Q\mathcal{G}$ gravity, where the Lagrangian density is a function of $Q$ and $\mathcal{G}$, $f(Q,\mathcal{G})$,
$S_{Q\mathcal{G}}=\int d^{4}x\sqrt{-g}f\left(Q,\mathcal{G}\right)\,.$ (21)
@@ IV Ghost problem
Let us remark that the Einstein-$f\left(\mathcal{G}\right)$ gravity includes ghosts in general. Nevertheless, some ghost-free models were proposed in Nojiri:2018ouv . In $f(Q,\tilde{R})$ or $\mathcal{G}$, $f(Q,\tilde{R},\mathcal{G})$, there could appear the ghost. As well known, the kinetic energy of the ghosts is unbounded below and there occur instabilities of the system in the classical theory. In the quantum theory, as in the Fadeev-Popov ghosts in the gauge theories Kugo:1979gm , the ghosts generate the negative norm states, which tells that negative probabilities appear and therefore the ghosts violate the Copenhagen interpretation of the quantum theory. The existence of the ghost tells that the model is physically inconsistent and therefore the problem of the ghosts is very important. If we find any ghosts in any model, we need to abandon the model or modify the model not to include ghosts. First, we briefly review the ghost problem in the Einstein-$f\left(\mathcal{G}\right)$ gravity.
The action of the Einstein-$f\left(\mathcal{G}\right)$ gravity Nojiri:2005jg ; Cognola:2006eg ; Leith:2007bu ; Li:2007jm ; Kofinas:2014owa ; Zhou:2009cy ; Oikonomou:2015qha is given by
$S=\int d^{4}x\sqrt{-g}\left(\frac{1}{2\kappa^{2}}{\tilde{R}}+f(\mathcal{G})+\mathcal{L}_{\mathrm{matter}}\right)\,.$ (25)
@@
$S=\int d^{4}x\sqrt{-g}\left(\frac{1}{2\kappa^{2}}{\tilde{R}}-\frac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi+h\left(\chi\right)\mathcal{G}-V\left(\chi\right)+\mathcal{L}_{\mathrm{matter}}\right)\,.$ (26)
@@
$S=\int d^{4}x\sqrt{-g}\left(\frac{1}{2\kappa^{2}}{\tilde{R}}+h\left(\chi\right)\mathcal{G}-V\left(\chi\right)+\mathcal{L}_{\mathrm{matter}}\right)\,.$ (27)
@@
$S=\int d^{4}x\sqrt{-g}\left\{\frac{1}{2\kappa^{2}}{\tilde{R}}+\lambda\left(\frac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi+\frac{\mu^{4}}{2}\right)-\frac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi+h\left(\chi\right)\mathcal{G}-V\left(\chi\right)+\mathcal{L}_{\mathrm{matter}}\right\}\,.$ (30)
@@
V FLRW spacetime in $f\left(Q,\mathcal{G}\right)$ gravity and cosmological reconstruction
@@ VI.4 Autonomous dynamical system
Here we have used the relation $\frac{d}{dt}=H\frac{d}{dN}$ between the derivatives with respect to the cosmological time $t$ and $e$-folding number $N$. For the dynamical analysis of $f(Q)$ gravity, see also Paliathanasis:2023nkb ; Paliathanasis:2023raj .
For the autonomous dynamical system, it is often used dimensionless quantities but because we have not specified $f\left(Q,\mathcal{G}\right)$, we use the dimensional quantities like $Q$ and $\mathcal{G}$. The fixed point defined by $\frac{d\mathcal{G}}{dN}=\frac{dQ}{dN}=0$ corresponds to de Sitter spacetime, where the Hubble rate $H$ is a constant $H=H_{0}$. Then the conditions for the fixed point are,
@@ VI.5 Gravitational wave
Note that $Q$ and $\mathcal{G}$ are invariant under the variation $g_{\mu\nu}\to g_{\mu\nu}+h_{\mu\nu}$ with (99). Although the full expression of the equation describing the gravitational wave is very complicated and lengthy, we are also now interested in the propagation speed of the gravitational wave. For the purpose, we only need to investigate the terms including the second-order derivative of $h_{\mu\nu}$. Then Eq. (III) tells
@@
In the FLRW spacetime, where $Q$ and $\mathcal{G}$ only depends on the cosmological time $t$, the condition (101) can be satisfied is $f_{\mathcal{G}}$ is a linear function of $t$, $f_{\mathcal{G}}=f_{0}+f_{1}t$. Here $f_{0}$ and $f_{1}$ are constants. The condition (101) cannot be, however, satisfied for the static and spherically symmetric spacetime as in the case of the Einstein-Gauss-Bonnet theories Nojiri:2023jtf . Therefore the existence of the Gauss-Bonnet term makes the situation highly non-trivial and the term is constrained by the observations.
There could be other kinds of fluctuations than gravitational waves of massless and spin two. Because the model includes the matter which is perfect fluid, there appears a sound wave, whose speed is given by ${c_{\mathrm{sound}}}^{2}=\frac{dp}{d\rho}=w$. Here $p$ and $\rho$ are pressure and the energy density of the matter and $w$ is the EoS parameter $p=w\rho$. In the case of the $f(R)$ gravity, there is a propagating scalar mode but in the case of the $f(Q)$ gravity, as shown in Capozziello:2024vix , the only propagating mode in the flat background is a massless and spin-two mode corresponding to the gravitational wave and there are no other kinds of propagating modes like scalar mode. This may suggest that there could not be other propagating modes than the graviton although there is an open question of the ghost. Even in the modified gravity theories, we rewrite the FLRW equations in the form of $\frac{3H^{2}}{\kappa^{2}}=\rho_{\mathrm{eff}}$, $-\frac{1}{\kappa^{2}}\left(2\dot{H}+3H^{2}\right)=p_{\mathrm{eff}}$. Here $\rho_{\mathrm{eff}}$ and $p_{\mathrm{eff}}$ are the effective energy density and the effective pressure. Then formally we may define the sound speed by ${c_{\mathrm{sound}}}^{2}=\frac{dp_{\mathrm{eff}}}{d\rho_{\mathrm{eff}}}=\frac{\frac{dp_{\mathrm{eff}}}{dt}}{\frac{d\rho_{\mathrm{eff}}}{dt}}$, for example. We should note, however, there is not the mode propagating with the speed of $c_{\mathrm{eff}}$ in the modified gravity theories.
@@ VII $QR\mathcal{G}$ gravity
As we mentioned, $f(Q,\tilde{R})$ gravity is identical to $f(Q,B)$ in Capozziello:2023vne or $f(Q,C)$ gravity in Gadbail:2023mvu . So far, we have discussed $f$ gravity in detail. In this section, we briefly discuss a generalised example, that is, $f(Q,\tilde{R},\mathcal{G})$ gravity in (19) for completeness.
There might look to be a special class of $f(Q,\tilde{R},\mathcal{G})$ gravity, that is, the Einstein-$f(Q,\mathcal{G})$ gravity, whose action is given by
$S_{\mathrm{Einstein-}f(Q,\mathcal{G})}=\int d^{4}s\sqrt{-g}\left(\frac{\tilde{R}}{2\kappa^{2}}+f(Q,\mathcal{G})\right)\,.$ (102)
Because the difference between the scalar curvature $\tilde{R}$ and $Q$ is a total derivative, this model is equivalent to the $f(Q,\mathcal{G})$ gravity, where the redefined $f(Q,\mathcal{G})$ is given by $f_{\mathrm{redefined}}(Q,\mathcal{G})=\frac{Q}{2\kappa^{2}}+f(Q,\mathcal{G})$.
@@
$\,-\tilde{R}_{\mu\nu}f_{\tilde{R}}+\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}f_{\tilde{R}}-g_{\mu\nu}\tilde{\nabla}^{2}f_{\tilde{R}}-2f_{\mathcal{G}}{\tilde{R}}{\tilde{R}}_{\mu\nu}+4f_{\mathcal{G}}{\tilde{R}}_{\mu\rho}{\tilde{R}}_{\nu}^{\ \rho}-2f_{\mathcal{G}}{\tilde{R}}_{\mu}^{\ \rho\sigma\tau}{\tilde{R}}_{\nu\rho\sigma\tau}$
@@
$S_{\mathrm{general}}=\int d^{4}x\sqrt{-g}f_{\mathrm{general}}\left(Q,g_{\mu\nu},{\tilde{R}}_{\mu\nu\rho\sigma}\right)\,.$ (104)
Here $g_{\mu\nu}$ expresses the metric which is not included in $Q$ and ${\tilde{R}}_{\mu\nu\rho\sigma}$ in the arguments. The equation corresponding to (VII) can be easily found by using the following formula of the variation with respect to the metric,
@@
Of course, $f\left(Q,\mathcal{G}\right)$ gravity and $f\left(Q,{\tilde{R}},\mathcal{G}\right)$ gravity are special cases of the general model (104). For example, $f\left(Q,R,\mathcal{G}\right)$ corresponds to the following $f_{\mathrm{general}}\left(Q,g_{\mu\nu},{\tilde{R}}_{\mu\nu\rho\sigma}\right)$,
$f_{\mathrm{general}}$ $\,\left(Q,g_{\mu\nu},{\tilde{R}}_{\mu\nu\rho\sigma}\right)$
$=$ $\,f\left(Q,{\tilde{R}}=g^{\mu\nu}g^{\rho\sigma}{\tilde{R}}_{\mu\rho\nu\sigma},\right.$
@@
×$\qquad\qquad\left.\times{\tilde{R}}_{\alpha\beta\gamma\delta}{\tilde{R}}_{\mu\nu\rho\sigma}\right)\,.$ (106)